# RHuman: Real-Time 3D Human Appearance Rendering from a Single Image

### 10 Dec 2023 v1

@@ Abstract
Reconstructing 3D human appearance from a single image is crucial for achieving holographic communication and immersive social experiences. However, this remains a challenge for existing methods, which typically rely on multi-camera setups or are limited to offline operations. In this paper, we propose RHuman, the first approach for real-time inference and rendering of photorealistic 3D human appearance from a single image. The core of our approach is to combine the strengths of implicit texture fields and explicit neural rendering with our novel representation, namely Z-map. Based on this, we present an end-to-end network that performs high-fidelity color reconstruction of visible areas and provides reliable color inference for occluded regions. To further enhance the 3D perception ability of our network, we leverage the Fourier occupancy field to reconstruct a detailed 3D geometry, which serves as a prior for the texture field generation and provides a sampling surface in the rendering stage. Experiments show that our end-to-end method achieves state-of-the-art performance on both synthetic data and challenging real-world images and even outperforms many offline methods. The project page is available for research purposes at http://cic.tju.edu.cn/faculty/likun/projects/R2Human.
1 Introduction
The rapid evolution of digital technology has notably impacted computer graphics and computer vision, particularly in 3D human appearance reconstruction. This advancement not only enhances visual experiences but also paves the way for practical applications in AR/VR, enabling interactive platforms like holographic communication and immersive social experiences. However, existing methods heavily rely on multi-camera setups [31, 19, 15, 20] or are constrained to offline operations [1, 7, 16], leaving real-time 3D human rendering from a single image as an unresolved challenge. In this paper, we aim to render the photorealistic 3D human appearance in real-time from a single image, as shown in Fig. 1.
@@ 1 Introduction
In this paper, we introduce RHuman, a real-time framework for rendering human appearance, which uniquely combines flow-based rendering techniques with an implicit 3D human geometry representation to synthesize novel view images. Our method only requires a single image as input and achieves both high-quality and real-time rendering performance. Moreover, our method can generate novel view renderings of humans in various clothing without individual-specific training. Unlike previous flow-based rendering methods that omit points occluded in the source view, we introduce an intermediate representation, namely the Z-map, during the rendering process, which collects the source view depth of the rendered points of the target view and forms them to a 2D map. It helps lift the 2D image feature into a 3D texture feature field, while maintaining compatibility with 2D neural rendering networks. This unique capability enables our network to learn data-driven 2D appearance knowledge of clothed human images, resulting in more accurate renderings. By preserving the features of occluded points and leveraging the Z-map, our method effectively resolves depth ambiguities. Additionally, we employ an efficient 3D object representation known as Fourier occupancy fields (FOF) [4], which explicitly represents a 3D object as a multi-channel image. It can serve both as a prior for texture field generation and as a sampling surface during the rendering stage, avoiding the high computational costs caused by dense sampling. Our method paves the way for the practical applications of AR/VR, which can be applied in holographic communication in the future.
@@
We present a novel system for high-quality, real-time synthesis of human novel view images with only a single RGB image input. To the best of our knowledge, this is the first system to restore the full-body appearance of a 3D human in real-time from a single image.
@@ 2.1 Monocular 3D Human Reconstruction
Human reconstruction has long been a concern in the domain of computer vision and graphics. Methods for predicting human bodies from a single image represent 3D shapes primarily by estimating the properties of points in 3D space. Some previous works [29, 21] directly predict the occupancy field of a given segment in space through regression networks, but such methods have high memory requirements, which limits the spatial resolution of shape estimation. Another type of methods involves using implicit function networks to remove resolution constraints. Saito et al. [16] proposed an implicit function based on pixel-aligned features to reconstruct a 3D human from a single image. PIFuHD [17] further introduces normal maps to improve geometric details. Subsequent methods [30, 25, 24] improve the robustness of the results by incorporating parametric SMPL [12] priors. Although these methods can achieve realistic results, they often require significant computational resources and are time-consuming. Feng et al. [4] proposed an efficient 3D geometric representation called Fourier Occupancy Field, which can represent a high-quality 3D geometry as a multi-channel image. It establishes a strong link between 3D geometry and 2D images, significantly reducing the computational requirements for reconstruction. However, one limitation is its inability to estimate texture during the geometry reconstruction process.
@@ 4 Method
Our goal is to build an end-to-end trainable framework that can achieve photorealistic novel view rendering of humans from a single RGB image. We employ FOF [4] to represent the 3D human form as a multi-channel image, thereby reducing computational time and memory overhead. Additionally, we introduce Z-map to effectively address the issue of prediction ambiguity for occluded points within the input image. Consequently, our network is capable of producing high-quality renderings of novel views using just a single image. As shown in Fig. 2, our framework mainly consists of three parts: 1) We use a reconstruction network to generate a Fourier occupation field from a single image. 2) We develop an image encoder to generate a texture field aligned with the reconstructed 3D geometry. 3) Finally, we warp the texture field and obtain other priors, and then use a rendering network to synthesize results with high visual quality.
@@ 4.1 Z-map
As illustrated in Fig. 3, the Z-map establishes a depth correspondence between the source and novel views. It aligns with the novel view; however, each pixel in the Z-map represents the depth of its corresponding point in the source view. Utilizing the source view’s depth information enables the decoder to more effectively leverage features from the input image. Additionally, the depth value in Z-map makes it possible to produce different colors for a specific image feature vector, which helps the network to perform more reasonable inference for invisible points.
@@ 4.2 RHuman Networks
4.2.1 Pixel-aligned Feature Encoder.
To leverage the geometric information obtained from the reconstruction network and enhance feature extraction, our encoder incorporates additional information derived from the reconstructed geometry. This allows for a more comprehensive understanding of the input image. This approach consists of three key components: FOF, normal map, and IUV map.
@@
IUV map. we add an auxiliary pixel-level dense correspondence as a prior knowledge to improve the reliability of network extraction because of the regularity of human textures. Specifically, we use the IUV map defined in DensePose [5] as the dense correspondence representation, which includes the partitioning and UV parametrization of the mesh surface.
@@
where is the concatenation operation. are the input color images, the predicted FOF, estimated surface normal and IUV maps, respectively. Note that our IUV map is rendered based on SMPL models [12] estimated from the input images.
4.2.2 Novel View Rendering.
Similar to the encoder, our rendering network also integrates some additional information to improve rendering performance. Note that we do not integrate Fourier occupation fields because rendering only focuses on the visible part of the image. We use novel camera parameters to render the normal map and IUV map, respectively. In addition to using the Z-map mentioned in Sec. 4.1, we also calculate soft masks as conditions to enable the network to understand the occlusion relationships between different views. We define the soft mask as the difference between the depth of the source image and the depth projection of the novel view onto the source view:
Then we warp the features of the source view to the novel view based on the flow map:
@@
where are the normal map and IUV map rendered by novel camera , respectively.
4.3 Loss Function
The loss functions used have a significant impact on the details recovered by the final model. For our rendering network, the output is a high-resolution image rendered from the target camera . We used the weighted sum for multiple loss terms as follows:
where is the common loss for direct supervised pixel points between and ground truth ; is the LPIPS perceptual loss [28] computed over pairs (); is 1 and is 0.5 in our experiments. As shown in the figure[], the LPIPS perceptual loss can effectively improve the reconstruction effect of the model on faces when the angle transformation is large.
@@ 5 Experiments
5.1 Implement Details
Dataset. Our our network is trained on synthetic data including pairs of meshes and rendered images. We collect 526 high-quality human scans from THuman2.0 dataset [27] with a wide range of clothing, poses and shapes. We randomly split them into a training set of 368 scans and a testing set of 105 scans. The remaining subjects are used as the validation set. For each scan, we render 32 views of high-resolution images with 16 lighting settings randomly selected from 630 lightings.
Training. For training our network, instead of using FOF constructed from the ground truth geometry, we apply the FOF-based reconstruction network [4] to predict the FOF representation from the single-view input image. Using the predicted FOF for training can make the rendering network obtain stronger generalization ability and improved robustness. We implement our method using PyTorch and train all network components jointly, end-to-end, using the Adam optimizer, with a learning-rate of . We train the network for 10 epochs which takes about 4 days with a batch size of 5 using a single NVidia RTX3090 GPU.
@@ 5.2 Comparisons
Qualitative results. Fig. 4 and Fig. 5 presents qualitative comparisons of our approach to the baseline on the 2k2k dataset [6] and the THuman2.0 dataset [27], respectively. PIFu demonstrates the ability to estimate reasonably accurate colors, but its performance is hindered by limitations in geometric estimation, resulting in suboptimal image quality. On the other hand, while SHERF heavily relies on SMPL priors, it struggles to synthesize satisfactory results for larger body sizes or those wearing loose-fitting clothing. Fig. 6 shows a qualitative comparison with traditional texture mapping methods. 3DTexture has a noticeable concatenation artifacts in the multi-view settings and cannot render occluded areas in the input view. Thanks to Z-map, Rhuman is able to incorporate geometric priors to enhance the rendering of occluded areas by 3D information. Furthermore, with the aid of neural rendering, Rhuman can effectively compensate for any inaccuracies in the geometric estimates. As a result, Rhuman is capable of synthesizing high-fidelity results that exhibit remarkable 3D consistency.
@@ 5.3 Ablation study
We are conducting ablation experiments to compare variants of our architecture and different training strategies.
@@
| zmap | normal | iuv | mask | SSIM | PSNR | LPIPS |
| ✓ | ✓ | ✓ | 0.9440 | 29.57 | 0.0405 | |
| ✓ | ✓ | ✓ | 0.9438 | 29.59 | 0.0402 | |
| ✓ | ✓ | ✓ | 0.9433 | 29.72 | 0.0395 | |
| ✓ | ✓ | ✓ | 0.9476 | 30.18 | 0.0394 | |
| w/o FOF | 0.9428 | 29.50 | 0.0401 | |||
| w/o VGG loss | 0.9423 | 29.37 | 0.0501 | |||
| Ours(full) | 0.9463 | 30.08 | 0.0392 | |||
Effects of additional information in the rendering network. Fig. 8 illustrates the comparison of the effects of incorporating additional information into the rendering network. This information is not generated by a neural network but obtained through computer graphics techniques. It is evident that the absence of each representation results in a noticeable decline in visual quality. Tab. 4 presents our quantitative results, with our full approach achieving the highest rank in terms of LPIPS, while ranking second in SSIM and PSNR. This outcome can be attributed to the fact that we do not utilize a ground truth mesh, but instead rely on estimated geometry to compute the mask. As a result, while our visual results appear accurate, there may be some small misalignment in terms of geometric consistency.
@@
Effects of perceptual loss. To explore the effects of perceived loss on training, we retrained a variant without using the perceived loss. Fig. 9 compares results obtained with and without the use of perceived loss, and we also make a quantitative comparison in the seventh line of Tab. 4. The results show that the perceptual loss allows the network to learn more texture details, thus improving the realism of the rendering results.
6 Results on 1K Resolution

### 21 Mar 2024 v2

@@ Abstract
Rendering 3D human appearance in different views is crucial for achieving holographic communication and immersive VR/AR. Existing methods either rely on multi-camera setups or have low-quality rendered images from a single image. In this paper, we propose RHuman, the first approach for real-time inference and rendering of photorealistic 3D human appearance from a single image. The core of our approach is to combine the strengths of implicit texture fields and explicit neural rendering with our novel representation, namely Z-map. Based on this, we present an end-to-end network that performs high-fidelity color reconstruction of visible areas and provides reliable color inference for occluded regions. To further enhance the 3D perception ability of our network, we leverage the Fourier occupancy field as a prior for generating the texture field and providing a sampling surface in the rendering stage. We also propose a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view coherence. Experimental results show that our method outperforms the state-of-the-art methods on both synthetic data and challenging real-world images, in real time. The source code will be available for research purposes.
keywords:
3D human appearance, rendering, single image, real-time
Given a single RGB image, RHuman can generate photorealistic 3D human appearance in real-time. We utilized the input (a) to render close views (b) and zoomed in (c). The results of pitch angle changes are shown in (d-e), and (f-h) demonstrate the outcomes of divergent views.
Introduction
The rapid evolution of digital technology has notably impacted computer graphics and computer vision, particularly in 3D human appearance reconstruction. This advancement not only enhances visual experiences but also paves the way for practical applications in AR/VR, enabling interactive platforms like holographic communication and immersive social experiences. However, existing methods heavily rely on multi-camera setups [34, 19, 15, 20] or are constrained to offline operations [1, 7, 16], leaving real-time 3D human rendering from a single image as an unresolved challenge. In this paper, we aim to render the photorealistic 3D human appearance in real-time from a single image, as shown in Fig.RHuman: Real-Time 3D Human Appearance Rendering from a Single Image.
@@ Introduction
In this paper, we introduce RHuman, a real-time framework for rendering human appearance from a single image, which uniquely combines flow-based rendering techniques with an implicit 3D human geometry representation to synthesize novel view images. Our method only requires a single image as input and achieves both high-quality and real-time rendering performance. Moreover, our method can generate novel view renderings of humans in various clothing without individual-specific training. Unlike previous flow-based rendering methods that omit points occluded in the source view, we introduce an intermediate representation, namely the Z-map, during the rendering process, which collects the source view depth of the rendered points of the target view and forms them to a 2D map. It helps lift the 2D image feature into a 3D texture feature field, while maintaining compatibility with 2D neural rendering networks. This unique capability enables our network to learn data-driven 2D appearance knowledge of clothed human images, resulting in more accurate renderings. By preserving the features of occluded points and leveraging the Z-map, our method effectively resolves depth ambiguities. Additionally, we employ an efficient 3D object representation known as Fourier occupancy fields (FOF) [4], which explicitly represents a 3D object as a multi-channel image. It can serve both as a prior for texture field generation and as a sampling surface during the rendering stage, avoiding the high computational costs caused by dense sampling. In order to ensure the consistency between multiple views and reduce the jittering phenomenon between neighboring frames, we propose a consistency loss to regularize this process and design a spatio-temporal fusion strategy to enhance this in practical applications. Our method paves the way for the practical applications of AR/VR, which can be applied in holographic communication in the future. Source code will be available for research purposes.
@@
We present a novel system for high-quality, real-time synthesis of human novel view images with only a single RGB image input. To the best of our knowledge, this is the first system to restore the full-body appearance of a 3D human in real time from a single image, paving the way for practical applications in AR/VR.
@@
We propose a consistency loss to ensure the multi-view coherence and reduce the jittering phenomenon, and design a spatio-temporal fusion strategy to enhance this in practical applications.
@@ 1.1 Monocular 3D Human Reconstruction
Human reconstruction has long been a concern in the domain of computer vision and graphics. Methods for predicting human bodies from a single image represent 3D shapes primarily by estimating the properties of points in 3D space. Some previous works [33, 22] directly predict the occupancy field of a given segment in space through regression networks, but such methods have high memory requirements, which limits the spatial resolution of shape estimation. Another type of methods involves using implicit function networks to remove resolution constraints. Saito et al. [16] proposed an implicit function based on pixel-aligned features to reconstruct a 3D human from a single image. PIFuHD [17] further introduces normal maps to improve geometric details. Subsequent methods [32, 27, 26, 30] improve the robustness of the results by incorporating parametric SMPL [12] priors. Although these methods can achieve realistic results, they often require significant computational resources and are time-consuming. Feng et al. [4, 5] proposed an efficient 3D geometric representation called Fourier Occupancy Field, which can represent a high-quality 3D geometry as a multi-channel image. It establishes a strong link between 3D geometry and 2D images, significantly reducing the computational requirements for reconstruction. However, one limitation is its inability to estimate texture during the geometry reconstruction process.
@@ 3 Method
Our goal is to build an end-to-end trainable framework that can achieve photorealistic novel view rendering of humans from a single RGB image. We employ FOF [4] to represent the 3D human form as a multi-channel image, thereby reducing computational time and memory overhead. Additionally, we introduce Z-map to effectively address the issue of prediction ambiguity for occluded points within the input image. Consequently, our network is capable of producing high-quality renderings of novel views using just a single image.
As shown in Fig. 1, our framework mainly consists of two parts: 1) We develop an image encoder to generate a texture field aligned with the reconstructed 3D geometry. 2) We warp the texture field and obtain other priors, and then use a rendering network to synthesize results with high visual quality. At inference, we estimate a detailed mesh from the input image, and using a pixel-aligned feature encoder to extract texture features. Next we mitigate depth blur in rendering with the help of Z-map representations, resulting in high-quality textures. During training, we use the input image to render two supervised views. We warp the first image to the second and perform visibility inference. By supervising the color of the visible points, we ensure the multi-view consistency of the results.
@@ 3.1 Z-map
3.1.1 Definition of Z-map
We propose a new representation called Z-map, which combines the strengths of flow-based rendering with implicit-field-based rendering. Unlike existing approaches that utilize the depth-map of the novel view for appearance rendering, our Z-map comprises the source view -coordinate of the visible points in the novel view, aligned to the novel view. By leveraging the source view image features and the -coordinate, the Z-map uniquely determines the position of visible points in the implicit field, thus eliminating depth ambiguity.
@@ 3.2 RHuman Networks
3.2.1 Pixel-aligned Feature Encoder
To leverage the geometric information obtained from the reconstruction network and enhance feature extraction, our encoder incorporates additional information derived from the reconstructed geometry. This allows for a more comprehensive understanding of the input image. This approach consists of two key components: FOF and normal map.
@@ 3.2.1 Pixel-aligned Feature Encoder
where is the concatenation operation. are the input color image, the predicted FOF and estimated surface normal map, respectively.
3.2.2 Novel View Rendering
Similar to the encoder, our rendering network also integrates some additional information to improve rendering performance. Note that we do not integrate Fourier occupation fields because rendering only focuses on the visible part of the image. We use novel camera parameters to render the normal map. Then we warp the features of the source view to the novel view based on the flow map:
@@ 3.2.2 Novel View Rendering
where is the normal map rendered by novel camera .
3.2.3 Spatio-Temporal Fusion Strategy
To maintain consistency during the rotation of the view, we introduce a spatio-temporal fusion strategy for generating the free-view video stream. Specifically, during the rotation process, we keep the image of the previous frame , and use Eq. 11 to get the surface points that are visible in both the current and previous frames. Subsequently, the color of in the current frame image is interpolated as:
where , is the rotation angle of the current view relative to the input view. The closer the current view is to the input view, the more we prefer the color of , while the farther the current view is from the input view, the more we use to ensure consistency. We set to in our experiments.
3.3 Training
We use synthetic data to train RHuman during the training stage. As illustrated in Fig. 1, an input image along with its corresponding FOF and normal map, sampled from the training set, serves as the input. The encoder then generates the corresponding texture field features, denoted as . With the camera parameters (comprising camera direction and position), we compute the warping flow which contains the novel view information we need, enabling the generation of an image of any view with the decoder. In each gradient step, we synthesize two images , with the same texture field features from a reference camera and an additional camera to perform multi-view consistency supervision.
Consistency Loss. To guarantee consistency between multiple views and reduce the jittering phenomenon during the rotation of the view, we propose a consistency loss to supervise the multi-view images rendered from the same texture field features. Specifically, we use Eq. 6 to calculate the flow map between the reference camera and the multi-view camera . Then we warp the image rendered by the reference camera to the muti-view camera view to get . Visibility inference is then performed: If the difference between the projection depth of reference view and the multi-view depth is lower than a threshold, we treat coordinate point v as visible:
where is a hyper-parameter and we set it to 0.02 in all our experiments.
We use the loss to supervise the visible points in the reference camera and the multi-view camera. The loss function can be expressed as:
where is the set of visible points between reference camera and multi-view camera calculated by Eq. 11.
Pixel Loss. Let the ground truth of the reference view and multi-view be and . We add an loss to supervise the pixel points between both sets of pairs (, ) and (, ), and to make the network more focused on the human, we only supervise the human foreground region of the image:
where is the set of foreground pixels of .
LPIPS Loss. In addition, to enhance the visual effect of the output image, we also add the LPIPS loss described in [31] for both sets of pairs (, ) and (, ):
where is the network described in [21].
Finally, our loss can be formulated as follows:
where we set , and to 100, 1 and 0.5 in our experiments.
Implement Details. Our our network is trained on synthetic data including pairs of meshes and rendered images. We collect 526 high-quality human scans from THuman2.0 dataset [29] with a wide range of clothing, poses and shapes. We randomly split them into a training set of 368 scans and a testing set of 105 scans. The remaining subjects are used as the validation set. For training our network, instead of using FOF constructed from the ground truth geometry, we apply the FOF-SMPL reconstruction network [4] to predict the FOF representation from the single-view input image. Using the predicted FOF for training can make the rendering network obtain stronger generalization ability and improved robustness. We implement our method using PyTorch and train all network components jointly, end-to-end, using the Adam optimizer, with a learning-rate of . We train the network for 10 epochs which takes about 4 days with a batch size of 4 using a single NVidia RTX3090 GPU.
@@ 4 Experiments
4.1 Real-time Rendering System
Our pipeline, designed for real-time monocular human novel view synthesis, operates as follows: Images are captured from the video stream using OpenCV [3] and preprocessed to a resolution of . We then apply RVM [11] for human mask extraction, effectively removing the background, and use HMR2.0 [8] for SMPL model estimation. Subsequently, the FOF-SMPL [4] is utilized to infer the occupation field. CUDA is employed to accelerate the rendering of depth, normal, and IUV maps. The final high-fidelity novel view images are rendered in real-time with the camera parameters and the previously processed data. Our pipeline, implemented with TensorRT on a single RTX-3090 GPU, achieves a performance of 24+ FPS, with the potential for further enhancement via additional GPUs. Fig. 2 showcase some of the real-time reconstruction results. Although our method trains to process each frame separately using synthetic data, it can still provide reasonable temporal consistency.
| & SMPL | information | ||||
| Time(ms) | 4.21 | 10.59 | 8.77 | 0.93 | 13.13 |
Tab. 2 details the execution time for each component of our pipeline. Currently, we implement our pipeline serially, achieving an overall runtime of approximately 37ms. We have combined the neural network and recovering of the occupancy field into a single component, optimized with TensorRT for enhanced acceleration. Moreover, there is potential for substantially increasing the pipeline’s processing speed through parallel implementation of each component.
@@ 4.2 Comparisons
Qualitative results. Fig. 3 and Fig. 4 presents qualitative comparisons of our approach to the baseline on the 2k2k dataset [6] and the THuman2.0 dataset [29], respectively. In addition, we tested the models on in-the-wild images, as shown in Fig. 5. PIFu demonstrates the ability to estimate reasonably accurate colors, but its performance is hindered by limitations in geometric estimation, resulting in suboptimal image quality. On the other hand, while SHERF can produce results with accurate pose, it encounters challenges in generating desirable outcomes for individuals wearing loose clothing, mainly due to its heavy reliance on SMPL priors. Additionally, since SHERF estimates colors for sampling points independently, it does not effectively consider the interrelationship between these points, often leading to ambiguous results. Fig. 6 shows a qualitative comparison with traditional texture mapping methods. 3DTexture has a noticeable concatenation artifacts in the multi-view settings and cannot render occluded areas in the input view. Thanks to Z-map, Rhuman is able to incorporate geometric priors to enhance the rendering of occluded areas by 3D information. Furthermore, with the aid of neural rendering, Rhuman can effectively compensate for any inaccuracies in the geometric estimates. As a result, Rhuman is capable of synthesizing high-fidelity results that exhibit remarkable 3D consistency.
@@
Comparison of running times. Tab. 5 shows the comparison results in terms of running time. We use TensorRT to accelerate inference in the real-time system. It can be seen that our method runs significantly faster than the other two baseline methods, with a speed improvement of two orders of magnitude.
| PIFu | SHERF | Ours | Ours (TensorRT) | |
| Time (ms) | 5741.62 | 1257.90 | 92.09 | 13.13 |
@@ 4.3 Ablation study
We conducted ablation experiments to compare variants of our architecture, different training strategies and different testing strategies.
@@
Effects of Z-map in the decoder. We retrained a variant without Z-map to explore the impact of Z-map. The first line of Tab. 6 shows the quantitative comparison results, and Fig. 7 shows the qualitative comparison results. It can be seen that Z-map can effectively improve the details of the generated results.
Effects of normal map in the rendering network. Fig. 8 illustrates the effects of incorporating normal maps into the rendering network. Normal maps are not generated by neural networks but are rendered from the reconstructed mesh. The second line of Tab. 6 shows the quantitative results. Normal maps can provide the network with more surface information, which allows the network to better obtain the correct color from the texture field and reduce the generation of artifacts.
Effects of FOF in the encoder. Fig. 10 shows the effect of using FOF in the encoder on the results, and the third line of Tab. 6 shows the quantitative results. FOF can help the encoder better understand the spatial information of the human mesh, thus enabling the network to synthesize accurate colors for large invisible areas, when the viewing angle changes substantially.
| SSIM | PSNR | LPIPS | |
| w/o | 0.9570 | 30.18 | 0.0392 |
| w/o normal | 0.9571 | 30.26 | 0.0390 |
| w/o FOF | 0.9567 | 29.93 | 0.0394 |
| w/o LPIPS loss | 0.9382 | 25.81 | 0.0808 |
| Ours(full) | 0.9571 | 30.52 | 0.0387 |
@@
Effects of LPIPS loss. To explore the effects of perceived loss on training, we retrained a variant without using the perceived loss. Fig. 11 compares results obtained with and without the use of perceived loss, and we also make a quantitative comparison in the fourth line of Tab. 6. The results show that the LPIPS loss greatly improves the network’s ability to learn the texture information, resulting in photo-realistic rendering results.
Effects of consistency strategy. We add a consistency loss and a spatio-temporal fusion strategy to ensure the multi-view consistency of the results. To explore the impact of the spatio-temporal fusion strategy on the results, we retrain a variant without the consistency loss and use the method described in [23] to draw EPI images to qualitatively evaluate the view consistency. Fig. 9 compares the results with and without the spatio-temporal fusion strategy, and it can be seen that our spatio-temporal fusion strategy effectively enhances the view consistency during rotation.
4.4 Results on 1K Resolution